OpenStudy (anonymous):

Help Linear equations... (2, -3) is a solution of the equation 3x - y = 3.

OpenStudy (anonymous):

True or False Question...

Parth (parthkohli):

Substitute x with 2 and y with -3. Check if it goes.

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow 3(2) - (-3) = 3 }\) \(\Large \color{MidnightBlue}{\Rightarrow 6 + 3 = 3 }\) Now, you can guess ^_^

OpenStudy (apoorvk):

Just substitute the value of 'x' and 'y' with 2 and -3 respectively from the coordinates, and check if it satisfies the equation - if it does, then it's a solution, which also means that the point lies on the line.

OpenStudy (anonymous):

So True...

Parth (parthkohli):

No, cesar!

Parth (parthkohli):

True in a world where 9 = 3

OpenStudy (apoorvk):

Why true? Is 6+3 = 3???

OpenStudy (anonymous):

My bad lol Thanks Guys